In a poll of

556

human resource​ professionals,

42.6%

said that body piercings and tattoos were big grooming red flags. Complete parts​ (a) through​ (d) below.

​a) Among the

556

human resource professionals who were​ surveyed, how many of them said that body piercings and tattoos were big grooming red​ flags?
Answer #1

GIVEN THAT

P = 42.6 % = 0.426

N = 556

Among the 556 human resource professionals who were​ surveyed, how many of them said that body piercings and tattoos were big grooming red​flag

EXPECTED VALUE = N * P = 556 * 0.426

= 236.856

= 237